# Odd-Parity Altermagnetism Originated from Orbital Orders

### v1

@@ Abstract
Figure 3: (a) In $\tau=\rho$ sectors, each valley hosts two spin-polarized massive Dirac cones separated by $\Delta E=2J$. The dashed (solid) lines refers to bands with a Chern number of $C_{s,\tau}=-1/2$ ($C_{s,\tau}=1/2$); the color denotes the spin polarization (red for spin-up, and blue for spin-down). (b) The spin-resolved Fermi surfaces are obtained at Fermi energy $\mu_{1}=-1.5t_{0}+J+1.4$ and $\mu_{2}=-1.5t_{0}-J+1.4$ when $t_{\perp}=0$. (c) Mid-gap helical edge states. $t_{\perp}=0.1$. Common parameters for (b-c): $\{t,t_{0},t_{d},\delta,J\}=\{1,0.1,0.5,0,2\}$. (d) Topological phase diagram as a function of interlayer coupling $t_{\perp}$ and exchange field $J$. The spin Chern number becomes $C^{s}=+1$ (1-1) when the chemical potential lying in the bulk energy gap at $E=-1.5t_{0}\pm J$. Parameters: $\{t,t_{0},t_{d},\delta\}=\{1,0,0.5,0\}$.
Altermagnetic quantum spin Hall insulators.–In our framework, both the $p$- and $f$-wave models can realize quantum spin Hall insulator phases. A simple intuition can be gained by considering the limit of vanishing interlayer coupling. In this limit, the top and bottom layers decouple and behave as two independent ferromagnetic monolayers that break time-reversal symmetry. When the system parameters fall into the topological regime, each monolayer becomes a Chern insulator. Because the two layers are related by an effective time-reversal symmetry, their Chern numbers are equal in magnitude but opposite in sign, yielding a zero total (charge) Chern number but a nonzero spin Chern number [85], and realizing a quantum spin Hall insulator.
To make the discussion concrete, we focus on the $f$-wave model (the $p$-wave case is presented in SM [81]). Setting $t_{\perp}=0$ renders the layer index $\rho_{z}$ a good quantum number. Expanding the Hamiltonian around the two valleys K and $\textbf{K}^{\prime}$ up to linear order in momentum $\boldsymbol{q}$, we obtain
@@
where $\tau=\pm 1$ labels the K and $\textbf{K}^{\prime}$ valleys, and $\rho=\rho_{z}=\pm 1$ denotes the top and bottom layers. The sectors with $\tau=-\rho$ are topologically trivial due to the vanishing linear dispersion. In contrast, in $\tau=\rho$ sectors, the low-energy spectrum at each valley consists of two spin-polarized massive Dirac cones separated in energy by a gap of $\Delta E=2J$, as schematically shown in Fig. 3(a). The Chern number of each Dirac cone is given by $\mathcal{C}_{s,\tau}=-{\rm sgn}(\tau t_{d})/2$ for $\delta<3|t_{d}|/2$. Because the effective time-reversal symmetry $\bar{E}||\mathcal{M}_{z}\mathcal{T}]$ relates K valley in top (bottom) to the $\textbf{K}^{\prime}$ valley in bottom (top) layer, Dirac cones with opposite spin polarization and opposite layer indices appear at the same energy. The resulting system therefore becomes a quantum spin Hall insulator, characterized by the spin Chern number $\mathcal{C}^{s}=\frac{\mathcal{C}_{\uparrow,\rho_{z}}-\mathcal{C}_{\downarrow,-\rho_{z}}}{2}$, and hosts a pair of helical edge states within each of the two bulk gaps located at $E=-\frac{3t_{0}}{2}\pm J$, as shown in Fig. 3(c).
Importantly, this topological character persists even in the presence of a finite interlayer coupling $t_{\perp}$. Although it hybridizes the two layers and removes $\rho_{z}$ as a good quantum number, the spin Chern number remains unchanged as long as the bulk energy gap at $E=-\frac{3t_{0}}{2}\pm J$ does not close. In practice, the nontrivial topology survives until $t_{\perp}$ becomes comparable in magnitude to the exchange field $J$ [see Fig. 3(d)]. We emphasize that this topologically nontrivial behavior is a generic feature of the odd-parity altermagnets obtained in our framework, and stands in sharp contrast to previously proposed noncollinear $p$-wave magnets [66, 67], which are typically topologically trivial.
@@
Discussions and conclusions.—Using symmetry principles as a guide, we demonstrate that odd-parity altermagnetism can emerge from the interplay of orbital order and collinear antiferromagnetism. This mechanism yields quantum spin Hall insulator phases when the exchange field dominates the interlayer coupling, a regime achievable in magnetic van der Waals (MvdW) compounds. In our construction, because the layer-flip operation does not depend on specific monolayer symmetries, any MvdW system with interlayer antiferromagnetism [88, 89, 90, 91, 92, 93] can in principle be converted to an odd-parity altermagnet via symmetry engineering. Magnetic Janus monolayers are a representative family of materials that naturally fulfill the required symmetry conditions, as their structure inherently lacks both inversion and mirror symmetries [94]. The defining characteristic of odd-parity altermagnets, i.e., odd-parity spin-splitting, can be directly detected in experiments using spin-resolved ARPES [28, 27, 29, 95].

### v2

@@ Abstract
Figure 3: (a) The $\tau\rho=1$ sectors host two spin-polarized massive Dirac cones separated by a gap $\Delta E=2J$ at each valley. The dashed (solid) lines refer to bands with a Chern number of $C_{s,\tau}=-1/2$ ($C_{s,\tau}=1/2$); the color represents the spin polarization. (b) The spin-resolved Fermi surfaces are obtained at Fermi energy $\mu_{1}=-1.5t_{0}+J+1.4$ and $\mu_{2}=-1.5t_{0}-J+1.4$ when $\{t_{0},t_{\perp},J\}=\{0.1,0,2\}$. (c) Mid-gap helical edge states. $\{t_{0},t_{\perp},J\}=\{0.1,0.1,2\}$. (d) Topological phase diagram as a function of interlayer coupling $t_{\perp}$ and exchange field $J$. $t_{0}=0$. The spin Chern number becomes $C^{s}=+1$ (1-1) when the chemical potential lying in the bulk energy gap between the two upper (lower) bands. Common parameters for (b-d): $\{t,t_{d},\delta\}=\{1,0.5,0\}$.
Altermagnetic quantum spin Hall insulators.–In our framework, both the $p$- and $f$-wave models can realize quantum spin Hall insulator phases. The intuition is clear in the limit of vanishing interlayer coupling: the two layers decouple and behave as independent ferromagnetic monolayers that break time-reversal symmetry. When parameters fall into the topological regime, each monolayer becomes a Chern insulator. Because the two layers are related by an effective time-reversal symmetry, their Chern numbers cancel, leving a zero total (charge) Chern number but a finite spin Chern number [85], thus realizing a quantum spin Hall insulator.
To make the discussion concrete, we focus on the $f$-wave model (the $p$-wave case is presented in SM [81]). Setting $t_{\perp}=0$ makes the layer index $\rho_{z}$ a good quantum number. Expanding the Hamiltonian around the valleys K and $\textbf{K}^{\prime}$ up to linear order in momentum $\boldsymbol{q}$, we obtain
@@
where $\tau=\pm 1$ labels K and $\textbf{K}^{\prime}$ valleys, and $\rho=\rho_{z}=\pm 1$ denotes the top and bottom layers. The $\tau\rho=-1$ sectors are topologically trivial due to the vanishing linear dispersion. In contrast, the $\tau\rho=1$ sectors host two spin-polarized massive Dirac cones separated by a gap $\Delta E=2J$, as schematically shown in Fig. 3(a). Each Dirac cone carries a Chern number $\mathcal{C}_{s,\tau}=-{\rm sgn}(\tau t_{d})/2$ for $\delta<3|t_{d}|/2$. Because the effective time-reversal symmetry $\bar{E}||\mathcal{M}_{z}\mathcal{T}]$ relates K valley in the top (bottom) layer to the $\textbf{K}^{\prime}$ valley in the bottom (top) layer, Dirac cones with opposite spin and layer indices appear at the same energy. The system therefore becomes a quantum spin Hall insulator, characterized by the spin Chern number $\mathcal{C}^{s}=\frac{\mathcal{C}_{\uparrow,\rho_{z}}-\mathcal{C}_{\downarrow,-\rho_{z}}}{2}$, and hosts a pair of helical edge states within each of the two bulk gaps located at $E=-\frac{3t_{0}}{2}\pm J$, as shown in Fig. 3(c).
Importantly, this topological character survives finite interlayer coupling $t_{\perp}$, provided that the two upper (lower) bands in each spin sector are separated by the gaps around $E=-\frac{3t_{0}}{2}\pm J$ [see Fig. 3(d)]. Although $t_{\perp}$ hybridizes the layers and removes $\rho_{z}$ as a good quantum number, the spin Chern number remains unchanged as long as these gaps stay open. In practice, this situation arises when the exchange field $J$ is comparable to, or larger than, the band width $W$ at $t_{0}=0$. For $J\sim W$, a small $t_{\perp}$ can open the band gaps that are closed at $t_{\perp}=0$, driving the system into a topological phase, while for larger $J$ the gap is already stablized by the exchange field. Once the system is nontrivial, further increasing $t_{\perp}$ eventually closes and reopens the gaps, rendering the system trivial. We emphasize that this topologically nontrivial behavior is a generic feature of odd-parity altermagnets in our framework, and stands in clear contrast to previously proposed noncollinear $p$-wave magnets [66, 67], which are typically trivial.
@@
Discussions and conclusions.—Using symmetry principles as a guide, we demonstrate that odd-parity altermagnetism can emerge from the interplay of orbital order and collinear antiferromagnetism. This mechanism yields quantum spin Hall insulator phases when the exchange field becomes comparable to the band width, a regime achievable in magnetic van der Waals (MvdW) compounds. In our construction, because the layer-flip operation does not depend on specific monolayer symmetries, any MvdW system with interlayer antiferromagnetism [88, 89, 90, 91, 92, 93] can in principle be converted to an odd-parity altermagnet via symmetry engineering. Magnetic Janus monolayers are a representative family of materials that naturally fulfill the required symmetry conditions, as their structure inherently lacks both inversion and mirror symmetries [94]. The defining characteristic of odd-parity altermagnets, i.e., odd-parity spin-splitting, can be directly detected in experiments using spin-resolved ARPES [28, 27, 29, 95].